3. Determine the pH for a 0.25 M NaOH(aq) solution. Show all calculations in order to receive full credit
Answer #1

Concentration of aq. NaOH = 0.25 M

NaOH is strong base and is mono basic in nature. Therefore,

[NaOH] = [OH-] = 0.25 M

Calculate the pOH of the solution

pOH = -log[OH-]

pOH = -log(0.25) = 0.602

Now calculate the pH of soltuionusing following formula

pH + pOH = 14

pH = 14 - pOH

pH = 14 - 0.602 = 13.4

pH of 0.25M solution is 13.4.